1. Which of the following media contains the largest amount of toluene, C6H5CH3 (show your calculations)? (a) 1 million gallons of water containing 5 ppm of toluene (b) 100 cubic meters of soils (bulk density = 1.8 g/cu. cm) with 5 ppm of toluene (c) Anemptywarehouse(30m×15m×10m)with5ppmVtolueneinair(P=1atmandT = 20 oC).
Answer #1

Molecular weight of toluene = (12*7)+8 = 92 g/mol

a) volume of water = 10^6 gallon

Assuming unit weight of water = 9.81 KN/m3 , thus weight of water = 1000000*0.003785411*9.81 = 37134.88 KN

Toluene = 5ppm thus mass of toluene =

(5/1000000)*(37134.88/9.81) = 0.018927 kg

Thus moles of toluene = 0.018927/(92/1000) = 3.7854 moles

b) for 100 m3 of soil , mass of toluene = (5/1000000)*(1.8*100) g = 0.0009 g

Thus moles of toluene = 0.0009/92 = 9.78*10^-6 moles

c) for empty warehouse, volume of air = 30*15*10 = 4500 m3

Thus volume of toluene = 0.0225 m3

Thus moles of toluene = 1*0.0225 /(8.20573*293.15)

= 9.35*10^-6 moles .

Thus largest amount of toluene is contained in water.